# Clustering MOOC Programming Solutions to Diversify Their Presentation to Students

### v1

@@ Abstract.
In many MOOCs, whenever a student completes a programming task, they can see previous solutions of other students to find potentially different ways of solving the problem and learn new coding constructs. However, a lot of MOOCs simply show the most recent solutions, disregarding their diversity or quality.
To solve this novel problem, we adapted the existing plagiarism detection tool JPlag to Python submissions on Hyperskill, a popular MOOC platform. However, due to the tool’s inner algorithm, it fully processed only 46 out of 867 studied tasks. Therefore, we developed our own tool called Rhubarb. This tool first standardizes solutions that are algorithmically the same, then calculates the structure-aware edit distance between them, and then applies clustering. Finally, it selects one example from each of the largest clusters, taking into account their code quality. Rhubarb was able to handle all 867 tasks successfully.
We compared approaches on a set of 59 tasks that both tools could process. Eight experts rated the selected solutions based on diversity, code quality, and usefulness. The default platform approach of selecting recent submissions received on average 3.12 out of 5, JPlag — 3.77, Rhubarb — 3.50. Since in the real MOOC, it is imperative to process everything, we created a system that uses JPlag on the 5.3% of tasks it fully processes and Rhubarb on the remaining 94.7%.
@@ 1. Introduction
Massive open online courses (MOOCs) are becoming more and more popular nowadays, serving the ever-growing need for accessible education (Oh et al., 2020). Programming MOOCs usually consist of theoretical and practical parts that allow students to learn coding concepts step by step (Staubitz et al., 2015). The correctness of practical coding tasks is usually validated by automated testing (assessment) systems (Staubitz et al., 2015). However, because of the large number of students, it is difficult to control other aspects of writing code (Kinash, 2013), e.g., code quality or specific ways the task can be solved. At the same time, these are very important for students’ development (Tigina et al., 2023; Keuning et al., 2019; Renkl, 2014; Brusilovsky, 2001).
One of the possible ways to learn new approaches to solve the same task is sharing knowledge among students (Kavitha and Ahmed, 2015; Liu et al., 2014; Sein-Echaluce et al., 2016). This approach is being actively studied and is even already implemented in several popular MOOCs, e.g., Codewars (cod, [n. d.]) or Hyperskill (jet, [n. d.]). These platforms allow students to see the solutions of other users after they successfully solve the given task. The main problem here is that these solutions might not be diverse enough, since the majority of them are very similar.
Thus, it is crucial to cut the number of possible solutions to show only different ones to the student. As far as we know, such a problem has not yet been actively studied from a student perspective, but a lot of tools were developed to solve similar problems from the perspective of a teacher in related research areas of plagiarism detection (Prechelt et al., 2000; Schleimer et al., 2003), feedback generation (Glassman et al., 2015; Gulwani et al., 2018), or the clustering of submissions (Effenberger and Pelánek, 2021). Specifically, some tools like JPlag (Prechelt et al., 2000) and MOSS (Schleimer et al., 2003) were developed to detect plagiarism in student submissions and can cluster them according to a plagiarism level.
In this work, we strived to solve this problem not only for research purposes, but also to improve this functionality for Python tasks on Hyperskill (jet, [n. d.]), a large MOOC platform where the current approach is to simply show the latest solutions. First of all, we tried to adapt the existing approaches to this task. We chose JPlag (Prechelt et al., 2000) due to its maturity and popularity, and adapted it to be able to run for our purposes. Like many other popular tools, the algorithm (Wise, 1993) that JPlag uses to compare code snippets cannot handle short programs. This is an especially important disadvantage for Python solutions, since they can be short even for complex tasks.
@@
The platform’s team provided us with a dataset of 867 Python tasks with different levels of complexity and 305,584 solutions in total. Of these tasks, JPlag was only able to fully process 46 (5.3%), with 434 more (50.1%) being processed partially (some solutions being skipped by the tool) and 387 (44.6%) not processed at all, including some complex tasks. On the other hand, Rhubarb successfully processed 100% of the data.
Finally, to compare different approaches, we used 59 tasks, the ones for which JPlag was able to process at least 90% of the solutions. For each of them, three groups of five solutions were generated — (1) with the default platform approach of selecting recent submissions, (2) with JPlag, and (3) with Rhubarb. Then, eight experts rated these groups from the standpoint of diversity, code quality, and usefulness, with an average of 2.3 experts per task. The resulting average scores are 3.12 out of 5 for the existing platform approach of simply selecting the most recent solutions, 3.77 for JPlag, and 3.50 for Rhubarb, so both tools performed better than the default approach from Hyperskill. Since in the real MOOC, processing all tasks and solutions is crucial, we implemented a system that employs the better-performing JPlag on the 5.3% of tasks it can fully process, and Rhubarb  on the remaining 94.7%.
You can find the source code of Rhubarb in our GitHub repository (Artser et al., 2024b). Separately, you can find a library with standardizing transformations for Python (Artser et al., 2024a). Finally, the supplementary materials for the paper are available on Zenodo (Artser et al., 2024c).
@@
  • Rhubarb, a clustering tool for creating diverse solutions in MOOCs, which handles all the studied platform’s tasks.
@@
  • An evaluation of Rhubarb with eight experts on 59 real tasks from Hyperskill, a large MOOC platform, showing that it outperforms the default platform approach.
@@ 2. Motivating example
In Figure 1, we can see five different solutions to the same task about number remainders and functional programming in Python, e.g., the map and filter functions. Let us imagine that some student submitted solution (a) and we want to show them some others.
Firstly, we notice that solution (b) is not just similar, it’s algorithmically the same. In this study, we define two code fragments to be algorithmically the same if they can be obtained from one another using some common transformations that do not affect the logic and can be applied without understanding the code, e.g., renaming variables, adding comments, etc. If these two solutions were to be brought to a normal form, they would be the same, thus we do not even need to consider solution (b).
Solution (c) is a different case. It looks even more similar to (a), to the degree where it may take some time to find the difference, and so it also probably should not be shown to the student. However, it is actually algorithmically different, since it skips using the list function twice. This does not matter in this particular context (map produces an iterator (pyt, [n. d.]) that works for this task), but this is not always the case, and this change could alter the logic in different contexts. To take such cases into account, it is necessary to cluster similar solutions, since transformations will not cover them.
On the other hand, solutions (d) and (e) are very different both compared to previous solutions and to each other. Solution (d) is an iterative approach, while solution (e) is an impressive but difficult to read one-liner. From the standpoint of diversity, however, they are relevant and interesting to be shown to the student.
This example demonstrates several key points that influence our study: (1) we can use code transformations to standardize code snippets and merge them (e.g., (a) and (b)); (2) some minor changes can not be considered by standardization, but should be handled by a clustering algorithm (e.g. (c)); (3) the shortness of the solution does not always correlate with task complexity (e.g., (e)) and short submissions should also be handled in our context.
@@ 3. Background
The topic of clustering student submissions is being actively studied for various purposes. Some studies are aimed at finding plagiarism in student solutions (Prechelt et al., 2000; Schleimer et al., 2003), while others aim to show their variety: to help teachers provide feedback on programming assignments (Gulwani et al., 2018; Glassman et al., 2015), or generate human-readable descriptions to help teachers identify different approaches (Effenberger and Pelánek, 2021).
@@
Another popular plagiarism detection tool is MOSS (Schleimer et al., 2003), a Web service that uses an adaptation of the greedy string tiling algorithm (Wise, 1993). The key difference is that the fragments of code marked as similar appear in no more than N submissions. MOSS also allows not to take into account the code that is directly repeated between solutions, e.g., the pre-written templates provided by the task’s creator. The tool supports many languages, however, the source code of the tool is closed and cannot be reused for research purposes.
OverCode (Glassman et al., 2015) helps teachers provide appropriate feedback to students at scale. This tool reformats the code using abstract syntax tree (AST) analysis, but only in terms of style, removing comments, and renaming variables; the comparison itself is still done by string-matching its lines. Moreover, OverCode was developed for Python 2, did not update since 2016, and fails on fragments with newer constructs, e.g., solution (e) from Figure 1.
@@
Overall, it can be seen that of the mentioned tools, only JPlag can be adapted to solve our problem for large enough solutions.
Figure 2. The general overview of Rhubarb.
@@ 4. JPlag
We ran JPlag from command line by giving it the path to input and output directories, as well as the settings of the clustering: its type (agglomerative hierarchical clustering with complete linkage) and a threshold. We used this type of clustering, because it is understandable and interpretable (Sembiring et al., 2011). The fine-tuning of the threshold is described in Section 7. JPlag calculates the similarity between each pair of solutions using the Dice coefficient (Dice, 1945), which measures how similar two sets of tokens are, and then clusters them using the provided threshold.
@@
However, JPlag cannot properly process all the submissions. Firstly, solutions are ignored if they have fewer than the specified number of particular tokens (12 for Python), and we cannot simply decrease this value because it is a base requirement for greedy string tiling algorithms. Secondly, only a limited number of tokens is supported that can be used for the clustering itself. As a result, for the example in Figure 1, JPlag was only able to connect solutions (a) and (b), only partially parsed solutions (c) and (d), and entirely skipped solution (e) because of its size. Overall, in the provided platform dataset, JPlag could process all solutions for only 5.3% of tasks. More detailed statistics about this are provided in Section 6.
@@ 5. Rhubarb
To overcome the shortcomings of JPlag so that all the submissions on the Hyperskill platform could be processed, we developed our own tool called Rhubarb. Its overview is presented in Figure 2; below we describe each stage in detail.
@@ 5.1. Standardization of Submissions
The library can be found separately on GitHub (Artser et al., 2024a). We developed this set of transformations according to an existing study (Rivers and Koedinger, 2017) that described standardizing transformations for constructs commonly found in student submissions.
We expect that the presented library can be useful for other potential tasks, e.g., hint generation, clone detection, or verifying program correctness without using test cases. The library is written in Kotlin and can be used directly or via an API.
After the transformations are applied, the standardized solutions are saved in a so-called solution graph (see (2) in Figure 2). At the standardization stage, the solution graph does not contain any edges. The vertices consist of a unique ID, standardized code, and a list of solution IDs that correspond to it. Although standardization in this form already constitutes clustering of code solutions, it only combines solutions with exactly the same semantics. Solutions with similar approaches (such as solution (c) from Figure 1) may still be in different groups, hence the need for subsequent clustering.
@@ 5.2. Measuring Distance
The distance measuring process involves turning the previously obtained solution graph into a complete weighted graph (see (3) in Figure 2). For each pair of standardized solutions, the distance is calculated and added as an edge with a weight. Consider the code snippets from Figure 1. Code snippets (a) and (b) can be considered as one node (a-b), since they merge during the standardization step. The distance between nodes (a-b) and (c) is $5656$, while the distance between (d) and (e) is $530530$ (see Figure 4).
@@ 5.3. Hierarchical Agglomerative Clustering
The result of this step is another complete weighted graph (see (4) in Figure 2). Its vertices are clusters containing a unique cluster ID and a list of standardized solutions. The edge weights of the cluster graph are the inter-cluster distances calculated during clustering. Each cluster is converted from a list of standardized solutions to a list of original student solutions from the platform and saved. An example of a simple graph that corresponds to the examples from Figure 1 is shown in Figure 4, with both inter-solution and inter-cluster distances. As we can see, code snippets (a), (b), and (c) represent the same cluster, and the submissions (d), and (e) build their own clusters.
@@ 5.4. Selecting Representative Solutions
Finally, we need to select representative examples to present them to the user. Similar to JPlag’s pipeline, we take the five largest clusters, however, instead of selecting the first solutions within them, we additionally sort them by code quality. Specifically, we take the grade provided by Hyperstyle (Birillo et al., 2022), a code quality tool that is used in the studied Hyperskill platform, sort the solutions in the cluster, and select the example from the ones with the highest grade. This allows us to additionally ensure that the provided examples are of good quality. In principle, any other sorting can be used in Rhubarb, according to the needs of a particular platform, e.g., code quality, date, size, etc.
@@ 5.5. Limitations
The developed tool has several limitations. First, standardization can take a long time, since we need to apply 12 transformations, and building and resolving a PSI tree is a complex process. We already implemented the ability to serialize the result of this step to reuse the standardized submissions, and we plan to add parallelization in future work. Next, to cluster student submissions, a complete weighted graph is built. This can be time- and memory-consuming for a large number of submissions. To overcome this drawback, we plan to support incremental graph recalculation in future work.
6. Dataset
To fine-tune and evaluate our approach, we used the data provided to us by the Hyperskill platform (jet, [n. d.]). The platform’s policy allows to use anonymized data for research purposes, and the data did not include any personal or identifying information. The platform has Hyperstyle (Birillo et al., 2022) as the embedded code style tool, so the code quality grade is already included in the data. The dataset represents all correct Python solutions submitted within $1616$ weeks, a total of 305,584 solutions for 867 different coding tasks. The data contains the ID and the code of each submission, its code quality, and a timestamp. The timestamp helps to simulate the default platform’s approach and extract only the several last solutions.

### v2

@@ Abstract.
In many MOOCs, whenever a student completes a programming task, they can see previous solutions of other students to find potentially different ways of solving the problem and to learn new coding constructs. However, a lot of MOOCs simply show the most recent solutions, disregarding their diversity or quality, and thus hindering the students’ opportunity to learn.
In this work, we explore this novel problem for the first time. To solve it, we adapted the existing plagiarism detection tool JPlag to Python submissions on Hyperskill, a popular MOOC platform. However, due to the tool’s inner algorithm, JPLag fully processed only 46 out of 867 studied tasks. Therefore, we developed our own tool called Rhubarb. This tool first standardizes solutions that are algorithmically the same, then calculates the structure-aware edit distance between them, and then applies clustering. Finally, it selects one example from each of the largest clusters, thus ensuring their diversity. Rhubarb was able to handle all 867 tasks successfully.
We compared different approaches on a set of 59 real-life tasks that both tools could process. Eight experts rated the selected solutions based on diversity, code quality, and usefulness. The default platform approach of simply selecting recent submissions received on average 3.12 out of 5, JPlag — 3.77, Rhubarb — 3.50. To ensure both quality and coverage, we created a system that combines both tools. We conclude our work by discussing the future of this new problem and the research needed to solve it better.
@@ 1. Introduction
Massive open online courses (MOOCs) are becoming more and more popular nowadays, serving the ever-growing need for accessible education (Oh et al., 2020). Programming MOOCs usually consist of theoretical and practical parts that allow students to learn coding concepts step by step (Staubitz et al., 2015). The correctness of practical coding tasks is typically validated by automated testing (assessment) systems (Staubitz et al., 2015), however, because of the large number of students, it is difficult to control other aspects of writing code (Kinash, 2013), e.g., code quality or specific ways the task can be solved. At the same time, these are very important for students’ development (Tigina et al., 2023; Keuning et al., 2019; Renkl, 2014; Brusilovsky, 2001).
In some cases, teachers may prepare several different approaches to solve every task, however, this is not always feasible. One other way of learning these different approaches is sharing knowledge among students (Kavitha and Ahmed, 2015; Liu et al., 2014; Sein-Echaluce et al., 2016). This is currently being actively studied and is even already implemented in several popular MOOCs, e.g., Codewars (cod, 2024) or Hyperskill (jet, 2024). These platforms allow students to see the solutions of other users after they solve the given task. The main problem here is that these solutions might not be diverse enough, since the majority of them are very similar.
For this reason, it is crucial to cut the number of possible solutions to show only different ones to the student. As far as we know, such a problem has not yet been actively studied from a student perspective, but a lot of tools were developed to solve similar problems from the perspective of a teacher in related research areas of plagiarism detection (Prechelt et al., 2000; Schleimer et al., 2003) and feedback generation (Glassman et al., 2015; Gulwani et al., 2018). Specifically, several tools like JPlag (Prechelt et al., 2000) and MOSS (Schleimer et al., 2003) were developed to detect plagiarism in student submissions and can cluster them according to a plagiarism level.
In this work, we strived to explore this novel problem not only for research purposes, but also to improve this functionality for Python tasks on Hyperskill (jet, 2024), a large MOOC platform where the current approach is to simply show the latest solutions. First of all, we tried to adapt the existing approaches to this task. We chose JPlag (Prechelt et al., 2000) due to its simplicity, maturity, and popularity, and adapted it to be able to run for our purposes in the clustering mode. Like in many other popular tools, the algorithm (Wise, 1993) that JPlag uses to compare code snippets cannot handle short programs. This is an especially important disadvantage for Python solutions, since they can be short even for complex tasks.
@@
The platform’s team provided us with a dataset of 867 Python tasks with different levels of complexity and a total of 305,584 solutions to them. Of these tasks, JPlag was only able to fully process 46 (5.3%), with 434 more (50.1%) being processed partially (some solutions being skipped by the tool) and 387 (44.6%) not processed at all, including some complex tasks. On the other hand, Rhubarb successfully processed 100% of the data.
Finally, for the pilot evaluation of different approaches, we used 59 tasks, the ones for which JPlag was able to process at least 90% of the solutions. For each of them, three groups of five solutions were generated — (1) with the default platform approach of selecting recent submissions, (2) with JPlag, and (3) with Rhubarb. Then, eight experts rated these groups on a Likert scale from the standpoint of diversity, code quality, and usefulness, with an average of 2.3 experts per task. The resulting average scores are 3.12 out of 5 for the existing platform approach of simply selecting the most recent solutions, 3.77 for JPlag, and 3.50 for Rhubarb, so both tools performed better than the default approach from Hyperskill. Since in the real MOOC, processing all solutions is crucial, and the difference in the quality between JPlag and our approach is not drastic, we implemented a system that employs JPlag on the 5.3% of tasks it can fully process, and Rhubarb  on the remaining 94.7%.
You can find the source code of Rhubarb in our GitHub repository (Artser et al., 2024b). Separately, you can find a library with standardizing transformations for Python (Artser et al., 2024a) that can be useful in other tasks. Finally, the supplementary materials for the paper are available on Zenodo (Artser et al., 2024c).
@@
  • Rhubarb, a tailored clustering tool for selecting diverse Python solutions for showing in MOOCs.
@@
  • A pilot evaluation of the concept and Rhubarb with eight experts on 59 real tasks from Hyperskill, a large MOOC platform, showing that it outperforms the default platform approach, as well as a discussion of the results together with the potential future of the proposed problem and the solution.
@@ 2. Motivating Example
[…]
Examples of different students’ approaches to solving the same task. Solutions (a) and (b) are algorithmically the same, while solution (c) is not, but still very similar to them. On the other hand, solutions (d) and (e) are rather different.
In Figure 1, we can see five different solutions to the same task about number remainders and functional programming in Python. Let us imagine that a student submitted solution (a) and we want to show them some other, different ones.
Firstly, we notice that solution (b) is not just similar, it’s algorithmically the same. In this study, we define two code fragments to be algorithmically the same if they can be obtained from one another using some common transformations that do not affect the logic and can be applied without understanding the code, e.g., renaming variables, adding comments, etc. If these two solutions were to be brought to a standard form, they would be the same, thus we do not even need to consider solution (b) for showing to the student.
Solution (c) is a different case. It looks even more similar to (a), to the degree where it may take some time to find the difference, and so it also probably should not be shown to the student. However, it is actually algorithmically different, since it skips using the list function twice. This does not matter in this particular context (map produces an iterator (pyt, 2024) that works for this task), but this is not always the case, and this change could alter the logic in different contexts. To take such cases into account, it is necessary to cluster similar solutions, since transformations will not cover them.
On the other hand, solutions (d) and (e) are very different both compared to previous solutions and to each other. Solution (d) uses an iterative approach, while solution (e) is an impressive but difficult to read one-liner. From the standpoint of diversity, however, they are relevant and interesting to be shown to the discussed student.
This example demonstrates several key points that influence our study: (1) we can use code transformations to standardize code snippets and merge them (e.g., (a) and (b)); (2) some minor changes can not be considered by standardization, but should be handled by a clustering algorithm (e.g. (c)); (3) the shortness of the solution does not always correlate with task complexity (e.g., (e)), and short submissions should also be handled in our context.
@@ 3. Background
The topic of clustering student submissions is being actively studied. Some studies are aimed at finding plagiarism in student solutions (Prechelt et al., 2000; Schleimer et al., 2003), while others aim to show their variety: to help teachers provide feedback on programming assignments (Gulwani et al., 2018; Glassman et al., 2015), or generate human-readable descriptions to help teachers identify different approaches (Effenberger and Pelánek, 2021).
@@
Another popular plagiarism detection tool is MOSS (Schleimer et al., 2003), a Web service that also uses an adaptation of the greedy string tiling algorithm (Wise, 1993). The key difference is that the fragments of code marked as similar appear in no more than N submissions. MOSS also allows for the exclusion of code that is directly repeated between solutions, such as the pre-written templates provided by the task’s creator. The tool supports many languages, however, the source code of the tool is closed and cannot be reused for research purposes, including ours.
OverCode (Glassman et al., 2015) helps teachers provide appropriate feedback to students at scale. This tool reformats the code using abstract syntax tree (AST) analysis, but only in terms of style, removing comments and renaming variables; the comparison itself is still done by string-matching its lines. OverCode was developed for Python 2, did not update since 2016, and fails on fragments with newer constructs, e.g., solution (e) from Figure 1.
@@
Overall, it can be seen that of the mentioned tools, only JPlag can be adapted to solve our problem at scale.
@@ 4. JPlag
We ran JPlag from command line by giving it the path to input and output directories, as well as the settings of the clustering: its type (agglomerative hierarchical clustering with complete linkage) and a threshold. We used this type of clustering, because it is understandable and interpretable (Sembiring et al., 2011). The fine-tuning of the threshold is described in Section 6.2. JPlag calculates the similarity between each pair of solutions using the Dice coefficient (Dice, 1945), which measures how similar two sets of tokens are, and then clusters them using the provided threshold.
@@
However, JPlag cannot properly process all the submissions. Firstly, solutions are ignored if they have fewer than the specified number of particular tokens (12 for Python), and we cannot simply decrease this value because it is a base requirement for greedy string tiling algorithms. Secondly, only a limited number of tokens is supported that can be used for the clustering itself. As a result, for the example in Figure 1, JPlag was only able to connect solutions (a) and (b), only partially parsed solutions (c) and (d), and entirely skipped solution (e) because of its size. Overall, in the provided platform dataset, JPlag could process all solutions for only 5.3% of tasks. More detailed statistics about this are provided in Section 6.1.
@@ 5. Rhubarb
To overcome the shortcomings of JPlag so that all the submissions on the Hyperskill platform could be processed, we developed our own tool called Rhubarb. Its overview is presented in Figure 2, below we describe each stage in detail.
@@ 5.1. Standardization of Submissions
Figure 2. The general overview of Rhubarb.
[…]
The general overview of Rhubarb.
@@
The library can be found separately on GitHub (Artser et al., 2024a). We developed this set of transformations according to an existing study (Rivers and Koedinger, 2017) that described standardizing transformations for constructs commonly found in student submissions. We expect that the library can be useful for other potential tasks, e.g., hint generation, clone detection, or verifying program correctness without using test cases. The library is written in Kotlin and can be used directly or via an API.
After the transformations are applied, the standardized solutions are saved in a so-called solution graph (see (2) in Figure 2). At the standardization stage, the solution graph does not contain any edges. The vertices consist of a unique ID, standardized code, and a list of solution IDs that correspond to it. Although standardization in this form already constitutes clustering of code solutions, it only combines solutions with exactly the same semantics. Solutions with similar approaches (e.g., (c) from Figure 1) may still be in different groups, hence the need for subsequent clustering.
@@ 5.2. Measuring Distance
The distance measuring process involves turning the previously obtained solution graph into a complete weighted graph (see (3) in Figure 2). For each pair of standardized solutions, the distance is calculated and added as an edge with a weight. Consider the code snippets from Figure 1. Code snippets (a) and (b) can be considered as one node (a-b), since they merge during the standardization step. The distance between nodes (a-b) and (c) is $5656$, while the distance between (d) and (e) is $530530$ (see Figure 4), indicating that (a-b) and (c) are more similar to each other.
@@ 5.3. Hierarchical Agglomerative Clustering
[…]
The graph with distances between the standardized solutions and clusters for solutions from Figure 1.
@@
The result of this step is another complete weighted graph (see (4) in Figure 2). Its vertices are clusters containing a unique cluster ID and a list of standardized solutions. The edge weights of the cluster graph are the inter-cluster distances calculated during clustering. Each cluster is converted from a list of standardized solutions to a list of original student solutions from the platform and saved. An example of a simple graph that corresponds to the examples from Figure 1 is shown in Figure 4, with both inter-solution and inter-cluster distances. As we can see, code snippets (a), (b), and (c) represent the same cluster, and the submissions (d), and (e) build their own clusters, just as we discussed.
@@ 5.4. Selecting Representative Solutions
Finally, we need to select representative examples to present them to the user. Similar to JPlag’s pipeline, we take the five largest clusters, however, instead of selecting the first solutions within them, we additionally sort them by code quality. Specifically, we take the code quality grade provided by the Hyperstyle tool (Birillo et al., 2022) that is used in the studied Hyperskill platform. We then sort the solutions in the cluster based on this grade, and select the example from the ones with the highest grade. This allows us to additionally ensure that the provided examples are of good quality. In principle, any other sorting can be used in Rhubarb, according to the needs of a particular platform, e.g., code quality, size, etc.
@@ 5.5. Limitations
The developed tool has several limitations. First, standardization can take a long time, since we need to apply 12 transformations, and building and resolving a PSI tree is a complex process. We already implemented the ability to serialize the result of this step to reuse the standardized submissions, and we plan to add parallelization in future work. Next, to cluster student submissions, a complete weighted graph is built. This can be time- and memory-consuming for a large number of submissions. To overcome this drawback, we plan to support incremental graph recalculation.
6. Evaluation
6.1. Dataset
To fine-tune and evaluate our approach, we used the data provided to us by the Hyperskill platform (jet, 2024). The platform’s policy allows to use anonymized data for research purposes, and the data did not include any personal or identifying information. The platform has Hyperstyle (Birillo et al., 2022) as the embedded code style tool, so the code quality grade is already included in the data. The dataset represents all correct Python solutions submitted within $1616$ weeks, a total of 305,584 solutions for 867 different coding tasks. The data contains the ID and the code of each submission, its code quality, and a timestamp. The timestamp helps to simulate the default platform’s approach and extract only the several last solutions.